The monthly salaries of the employees at Lyman Limited follow a normal distribution with an average salary of R 14 000 and a standard deviation of R 500. ,If a female employee claims to be in the top 10% of highest paid employees at the company, what must be her minimum monthly salary


1
Expert's answer
2021-09-21T02:38:23-0400

"\mu = 14000 \\

\sigma= 500 \\

Z= \frac{x- \mu}{\sigma}"

For the top 10% P=0.90

From Z table:

"1.28 = \frac{x-14000}{500} \\

x = 14000 + (1.28 \times 500) \\

= 14640"